OpenStudy (anonymous):

if f(x)=cos^3(4x) then f'(x)=?

OpenStudy (tkhunny):

You have some chain rule in your future. Give it a go!

OpenStudy (anonymous):

(cos^3)d/dx(4x) + (4x) d/dx cos^3

OpenStudy (anonymous):

is that right?

OpenStudy (tkhunny):

?? There is no multiplication. \(f(x)=cos^{3}(4x)\) \(f'(x) = 3\cdot\cos^{2}(4x)\cdot (SOMETHING) \cdot (SOMETHING \; ELSE)\) Fill in the blanks.